# Hybrid Rendering for Multimodal Autonomous Driving: Merging Neural and Physics-Based Simulation

### v1

@@ 3.1.2 LiDAR intensity and depth map generation
To achieve this, we generate dense depth and intensity maps from LiDAR data. We first create a static aggregated point cloud by removing LiDAR points that fall within the bounding boxes of dynamic objects in each sweep, then merge the sweeps using per-point egomotion correction. Next, we generate sparse depth and intensity maps by projecting the point cloud onto the camera frames. We refine the depth maps using a classical depth-completion algorithm [13] and apply a simple depth-based dilation to the intensity maps.
@@ 4 Results and Experiments
4.1 Reconstruction quality
@@ 4.2 Downstream tasks
Semantic segmentation
@@
Figure 6: Comparison of learned segmentation with the Mask2Former model. Green regions correspond to matches between the two models, while blue and orange colors represent regions classified by only the learned segmentation or Mask2Former respectively.
Thirdly, we conducted a qualitative evaluation of our extreme novel view renderings of the road surface by employing a segmentation network to detect lanes, roadsides, and the road itself. As depicted in Figure 7, the network successfully detected these critical features in images rendered from a 5-meter height, laterally shifted by 2 meter, and with the camera pitched down by 25°.
@@
IoU3D Scenario AP3D [%] $\uparrow$ APH3D [%] $\uparrow$
All 0-30 30-50 50$\infty$ All 0-30 30-50 50$\infty$
0.3 Original 11.08 20.16 23.42 0.54 10.99 20.04 23.00 0.53
Shifted (0 m) 9.40 23.88 13.73 0.39 9.34 23.73 13.55 0.38
Shifted (1 m) 8.17 22.57 11.46 0.00 8.08 22.46 10.91 0.00
Shifted (2 m) 9.17 22.38 15.04 0.00 9.02 21.76 14.86 0.00
Shifted (3 m) 9.93 22.69 14.75 0.00 9.86 22.48 14.56 0.00
3D object detection.
Table 4 presents the detection results using Waymo detection metrics, where the ’Original’ scenario refers to using the original images, while ’Shifted’ scenarios are generated by our reconstruction model with laterally offset ego-trajectories. Interestingly, the model performs even better in novel view scenarios than on the original images at close range (0–30m), primarily due to a higher number of false positives in the original images. Distant object detection performance declines in novel view synthesis, warranting further investigation on a larger sample. However, detection results remain largely consistent across novel view scenarios. The observed differences can be attributed to object visibility changes caused by lateral shifts, as the model struggles to recognize truncated objects.
Qualitative results show that the model successfully detects vehicles rendered by both the reconstruction model and the mesh-based rendering engine, indicating that no significant domain gap is introduced. However, distant objects remain undetected due to the model’s detection range limitation. Example detections are shown in Figure 8. Additional figures showing the model’s detections on data rendered by our hybrid solution can be found in the Supplementary Material.
5 Limitations and Future Work
Limitations.
Future work.
Acknowledgment
Supplementary Material
Appendix A Qualitative results
A.1 Nerf-based reconstruction
@@ A.3 Hybrid rendering
A.4 Downstream task: Semantic Segmentation
Figure 9 illustrates a comparison between our learned segmentations and Mask2Former’s results for the ’CAR’ label. Regions marked in blue are classified as ’CAR’ by our learned segmentation model, but not by Mask2Former. Our approach achieves accurate segmentations, even for partially obstructed or distant objects, due to the association of class labels with discrete 3D object representations (Gaussians). These representations are trained from diverse viewpoints and distances, rather than relying solely on 2D pixel information. Consequently, this example demonstrates that the comparison metric presented in the main text may underestimate our actual performance, as it incorporates Mask2Former’s erroneous predictions.
A.5 Downstream task: 3D Object Detection
Figure 10 and 11 show detections from DEVIANT [14], a publicly available monocular model, on scenarios generated using our method. The reconstruction model rendered the static environment, while the mesh-based rendering engine introduced dynamic vehicles. As shown, the model successfully detects vehicles from both rendering methods, suggesting that no significant domain gap is introduced. Distant objects are not recognized due to model limitations.
@@ A.5 Downstream task: 3D Object Detection
Figures 13 and 12 show 3D bounding boxes detected by the DEVIANT model on the same frame after these horizontal shifts. As observed, viewpoint changes can help reduce false positives by improving object visibility. However, they can also introduce false negatives when certain objects become less visible after the shift. These effects help explain the metric differences presented in the paper.

### v2

@@ 3.1.2 LiDAR intensity and depth map generation
To achieve this, we generate dense depth and intensity maps from LiDAR data. We first create a static aggregated point cloud by removing LiDAR points that fall within the bounding boxes of dynamic objects in each sweep, Then merge the sweeps using per-point egomotion correction. Then we split the aggregated point cloud into square chunks in BEV space, and employ Poisson surface reconstruction [12] for each chunk to generate a dense, mesh-based representation of the lidar data. Finally, we use Open3D [40] to render the dense depth maps and LiDAR intensity images from these meshes using the intrinsic and extrinsic camera parameters of the train camera frames.
@@ 3.3 Rendering pipeline
Figure 4: Qualitative comparison of novel view synthesis results between OmniRe (left) and our model (right) for the static components of the scene. Differences in reconstruction quality are particularly noticeable in the rendering of lane boundaries and road markings.
@@ 4 Results and Experiments
4.1 Static scene reconstruction quality
@@
4.2 Novel view synthesis quality
We evaluate the novel view synthesis performance of our method in comparison with OmniRe [5], which is considered the state-of-the-art approach for supervised dynamic scene reconstruction. For quantitative evaluation, both OmniRe and our model were trained on the same scenes from the Waymo dataset using all five cameras, and a separate set of held-out images was used for metric computation. Because our method focuses on reconstructing only the static components of the environment, we report PSNR and SSIM scores computed exclusively over the static regions. Table 2 presents a comparison of our method and OmniRe across three dynamic Waymo scenes using these metrics. Additional implementation details are provided in the Appendix.
Although evaluating ”left out” images from the original dataset yields quantitative insights, this setup primarily measures the models’ interpolation capabilities and does not adequately reflect their robustness to significant viewpoint shifts relative to the training trajectory (e.g., lane changes), which are common in autonomous driving scenarios. To address this limitation, we also perform qualitative novel view synthesis comparisons between OmniRe and our method under laterally shifted ego trajectories of up to 5 meters.
Figures 4 and 7 demonstrate the qualitative differences in highly extrapolated views, focusing on autonomous driving critical features like lane markings, traffic signs, and pedestrian crossings. We assume our model achieves superior quality in these scenarios due to the explicit normal regularization. This regularization constrains the flat splats to align with scene surfaces, leading to a more realistic reconstruction of the actual geometry, in addition to the accurate rendering of RGB images.
Figure 7: Qualitative comparison of novel view synthesis quality between OmniRe (left) and our model (right).
@@ 4.3 Downstream tasks
[…]
13299463… 26.15/0.8011 26.07/0.7908
17860546… 30.13/0.9017 29.89/0.8894
30154365… 28.36/0.8767 27.62/0.8433
Table 2: Novel view synthesis quality comparison on Waymo scenes.
IoU3D Scenario AP3D [%] $\uparrow$ APH3D [%] $\uparrow$
All 0-30 30-50 50-$\infty$ All 0-30 30-50 50-$\infty$
0.3 Original 11.08 20.16 23.42 0.54 10.99 20.04 23.00 0.53
Shifted (0 m) 9.40 23.88 13.73 0.39 9.34 23.73 13.55 0.38
Shifted (1 m) 8.17 22.57 11.46 0.00 8.08 22.46 10.91 0.00
Shifted (2 m) 9.17 22.38 15.04 0.00 9.02 21.76 14.86 0.00
Shifted (3 m) 9.93 22.69 14.75 0.00 9.86 22.48 14.56 0.00
IoU3D AP3D [%] $\uparrow$ APH3D [%] $\uparrow$ Recall@0.95
0–30 30–50 0–30 30–50 0–30 30–50
0.5 100.00 52.94 99.05 52.09 100.00 0.00
0.7 66.67 40.00 65.92 39.87 0.00 0.00
Table 4: Comparison of 3D object detection results using the Detection Agreement metric.
3D object detection.
Table 3 presents the detection results using Waymo detection metrics, where the ’Original’ scenario refers to using the original images, while ’Shifted’ scenarios are generated by our reconstruction model with laterally offset ego-trajectories. Interestingly, the model performs even better in novel view scenarios than on the original images at close range (0–30m), primarily due to a higher number of false positives in the original images. Distant object detection performance declines in novel view synthesis, warranting further investigation on a larger sample. However, detection results remain largely consistent across novel view scenarios. The observed differences can be attributed to object visibility changes caused by lateral shifts, as the model struggles to recognize truncated objects. In addition, we evaluated the detection results on the original and simulated images using the Detection Agreement metric [15] (see Table 4). The strong agreement between detections on original and simulated images suggests a small domain gap.
Qualitative results show that the model successfully detects vehicles rendered by both the reconstruction model and the mesh-based rendering engine, indicating that no significant domain gap is introduced. However, distant objects remain undetected due to the model’s detection range limitation. Example detections are shown in Figure 8.
Additional figures showing the model’s detections on data rendered by our hybrid solution, as well as quantitative results of another downstream task (semantic segmentation), can be found in the Supplementary Material.
5 Limitations and Future Work
Limitations.
Future work.
Acknowledgment
Supplementary Material
Appendix A Quantitative results
A.1 Novel view synthesis quality comparisons
In addition to comparing against OmniRe [5], we further evaluated the novel view synthesis performance of our method relative to DeSiRe-GS [22], a state-of-the-art unsupervised dynamic scene reconstruction approach for urban environments. Table 5 reports the PSNR and SSIM scores for DeSiRe-GS alongside those of OmniRe and our method. In the case of OmniRe and DeSiRe-GS, we used their official repositories for training and metrics calculations. Across the evaluated scenarios, our approach matches or exceeds the performance of existing state-of-the-art neural reconstruction techniques. Note that our method uses a bilateral guided ISP disentanglement method. While this approach mitigates camera appearance inconsistencies (which are highly relevant in real-world applications), it also has a negative impact on our PSNR and SSIM scores.
Sequence Our 3DGS OmniRe DeSiRe-GS
[…]
13299463… 26.15/0.8011 26.07/0.7908 26.94/0.8071
17860546… 30.13/0.9017 29.89/0.8894 28.58/0.8499
30154365… 28.36/0.8767 27.62/0.8433 26.03/0.7908
Table 5: Novel view synthesis quality comparison on Waymo scenes. The highest-performing method is highlighted in bold, while the second-best result is indicated in italics.
A.2 Downstream task: Semantic Segmentation
Our evaluation leverages semantic segmentation to analyze model performance in three key areas. Firstly, we evaluate the accuracy of our 3D reconstructions. This is achieved by comparing segmentation masks produced by Mask2Former [6] on ground truth images with those generated from renderings of our NeRF and 3DGS models. To ensure a fair comparison, ground truth images are reprojected to align with the undistorted pinhole camera parameters used for rendering. We quantify the reconstruction quality by calculating the percentage of pixels with matching classifications. To mitigate potential bias from easily segmented regions like road surfaces and sky, we also report the Intersection over Union (IoU) specifically for the ’CAR’ class. It is important to note that this metric also incorporates errors from the Mask2Former model, for example in cases where our segmentation accurately predicts the labels for distant/obstructed objects but Mask2Former fails to do so, based on the more limited information.
@@ A.2 Downstream task: Semantic Segmentation
Figure 9: Comparison of learned segmentation with the Mask2Former model. Green regions correspond to matches between the two models, while blue and orange colors represent regions classified by only the learned segmentation or Mask2Former, respectively.
@@
Appendix B Qualitative results
B.1 Nerf-based reconstruction
@@ B.3 Hybrid rendering
B.4 Extreme novel view rendering
On Figure 11, we illustrate the typical challenges of vanilla 3D Gaussian Splatting when it comes to untextured surfaces and handling extremely novel views. The resulting depth maps exhibit large errors because using only the RGB images does not contain sufficient information for accurate geometry reconstruction, and without direct regularization, the normals of the Gaussians are ill-defined. This failure of the base method significantly hinders its usefulness in autonomous driving simulation and motivates the need for the specialized approaches evaluated in the following section. Figure 18 shows novel view renders generated from poses laterally shifted by up to $5\text{m}$ from the original trajectory, comparing the performance of DeSiRe-GS, OmniRe, and the method presented herein. Since ground-truth imagery is unavailable for these substantial spatial shifts, quantitative error metrics cannot be computed. The comparison thus relies on a qualitative analysis centered on features deemed crucial for autonomous vehicle perception, including lane markings, pedestrian crossings, traffic signs, and vehicle shapes. Across the examined Waymo scenes, DeSiRe-GS frequently generates visually unacceptable outputs characterized by significant blurring and floaters. Both OmniRe and our model, however, produce coherent renderings. The performance divergence between OmniRe and our approach is most pronounced when examining these perception-critical elements. Specifically, OmniRe introduces distorted lane markings and crosswalks and blurry traffic signs. In contrast, our proposed model consistently maintains superior fidelity and geometric integrity of these essential road features. For a dynamic assessment of these results, we refer the reader to the supplementary comparison videos.
Figure 11: RGB, depth, and normal images from our (left) and vanilla GS (right) models from extreme novel viewpoint
B.5 Downstream task: Semantic Segmentation
Figure 13 illustrates a comparison between our learned segmentations and Mask2Former’s results for the ’CAR’ label. Regions marked in blue are classified as ’CAR’ by our learned segmentation model, but not by Mask2Former. Our approach achieves accurate segmentations, even for partially obstructed or distant objects, due to the association of class labels with discrete 3D object representations (Gaussians). These representations are trained from diverse viewpoints and distances, rather than relying solely on 2D pixel information. Consequently, this example demonstrates that the comparison metric presented in the main text may underestimate our actual performance, as it incorporates Mask2Former’s erroneous predictions. We also conducted a qualitative evaluation of our extreme novel view renderings of the road surface by employing a segmentation network to detect lanes, roadsides, and the road itself. As depicted in Figure 10, the network successfully detected these critical features in images rendered from a 5-meter height, laterally shifted by 2 meter, and with the camera pitched down by 25°.
B.6 Downstream task: 3D Object Detection
Figure 14 and 15 show detections from DEVIANT [14], a publicly available monocular model, on scenarios generated using our method. The reconstruction model rendered the static environment, while the mesh-based rendering engine introduced dynamic vehicles. As shown, the model successfully detects vehicles from both rendering methods, suggesting that no significant domain gap is introduced. Distant objects are not recognized due to model limitations.
@@ B.6 Downstream task: 3D Object Detection
Figures 17 and 16 show 3D bounding boxes detected by the DEVIANT model on the same frame after these horizontal shifts. As observed, viewpoint changes can help reduce false positives by improving object visibility. However, they can also introduce false negatives when certain objects become less visible after the shift. These effects help explain the metric differences presented in the paper.
Appendix C System Performance and Limitations
C.1 Complexity
While the pipeline involves multiple stages, each component is modular, well-isolated, and produces human-interpretable outputs. This modularity allows for effective debugging, maintenance, and integration, as each step can be validated independently, reducing the likelihood of error propagation. Assuming valid input data, the stages are robust and not inherently error-prone. Furthermore, the simulation tool is already in use by industrial partners as a closed-source solution and is available for academic use under a free license. This demonstrates the system’s practicality and usability beyond a purely research context.
C.2 Block Boundary Artifacts
The difference in distant textures arose from a NeRF distance cutoff during initial point cloud generation. Removing this cutoff has significantly reduced block edge artifacts (seen in Fig. 12). To ensure a consistent sky, we propose training a single sky model from all images and freezing it during block training.
Figure 12: Images rendered from the same viewpoint from different blocks close to the block edge
C.3 Reconstruction efficiency/rendering speed
Our method takes 12 ms to render a Full HD image, and a 64-beam LiDAR frame takes 17 ms on an NVIDIA RTX 3090.
@@ C.3 Reconstruction efficiency/rendering speed
(a) Novel view images shifted by 3 meters from the ego trajectory on the Waymo segment-132… scene
(b) Novel view images shifted by 5 meters from the ego trajectory on the Waymo segment-178… scene
(c) Novel view images shifted by 4 meters from the ego trajectory on the Waymo segment-301… scene
(d) Novel view images shifted by 4 meters from the ego trajectory on the Waymo segment-301… scene
Figure 18: Extreme novel view renders with our model, OmniRe, and DeSiRe-GS, respectively